OpenStudy (anonymous):

Three persons Amar, Akbar and Anthony invested different amounts in a fixed deposit scheme for one year at the rate of 12% per annum and earned a total interest of Rs.3240 at the end of the year. If the amount invested by Akbar is Rs.5000 more than the amount invested by Amar and the amount invested by Anthony is Rs.2000 more than the amount invested by Akbar, what is the amount invested by Akbar?

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

x(1+12/100)-x=3240 but there are three investersx=27000

OpenStudy (anonymous):

amar=x,akbar=x+5000,anthony=x+7000

OpenStudy (anonymous):

12x/100=3240

OpenStudy (geekfromthefutur):

Is it multiple choice?

OpenStudy (anonymous):

r u asking multiple choice question

OpenStudy (geekfromthefutur):

Yes

OpenStudy (anonymous):

yes

OpenStudy (geekfromthefutur):

Rs 10000...

OpenStudy (geekfromthefutur):

Is that one of them cuz thats the correct answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@geekfromthefutur explain it

OpenStudy (anonymous):

3x+12000(1+12/100)-x=3240

OpenStudy (anonymous):

x=4513.

OpenStudy (geekfromthefutur):

1^7+2^7+3^7+4^7. 1000^7=x find remainder of x wid

OpenStudy (anonymous):

yeah x=5000

OpenStudy (anonymous):

see total x+x+5000+x+7000(1+12/100)-x=3240

OpenStudy (geekfromthefutur):

T

OpenStudy (geekfromthefutur):

Then thats ur answer

OpenStudy (anonymous):

whut i dont understand.1^7+2^7+3^7

OpenStudy (anonymous):

@ikram002p isit right